### v2

@@ 6.3 Interpolation from the FNN
Combining Theorem 6.1 and Theorem 6.4 leads to a next-token prediction capacity lower bound for one-layer multi-head decoder-only transformers.
Theorem 6.5.
Let $d,m_{0},d_{0},m,\omega\in\mathbb{N}$ and $d_{r}\in\mathbb{N}\leavevmode\nobreak\ \forall r\in[m_{0}]$. Let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Let $h_{\theta}$ be the one-layer multi-head decoder-only transformer model defined in Equation (LABEL:eq:model). Then the next-token prediction capacity of $\{h_{\theta}\mid\theta\}$ is greater than or equal to $m$.
@@
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\omega]^{t}$ be distinct and define $T=\max_{i\in[n]}|\alpha^{i}|$. Invoking Theorem 6.1, choose $(z,u)$ such that $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Invoking Theorem 6.4, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y^{\prime}\in(\mathrm{ri}\Delta^{\omega-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\omega\times n}$ such that $\phi(Y)=Y^{\prime}$ where $\phi$ is the column-wise softmax function. Furthermore, there exists $V\in\mathbb{R}^{m\times\omega}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$. So, by defining $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d},U=\mathbbm{1}_{d}u^{\top}/\sqrt{d},W_{1,r}=W_{2,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{r}}^{\top}/\sqrt{dd_{r}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{3,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{0}}^{\top}/\sqrt{dd_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}=\mathbbm{1}_{m_{0}d_{0}}\mathbbm{1}_{d}^{\top}/(m_{0}\sqrt{dd_{0}}),W=\mathbbm{1}_{d}w^{\top}/\sqrt{d},b=\eta\mathbbm{1}_{m}$, we get that $Y^{\prime}=[h_{\theta}(\alpha^{i})]_{i\in[n]}$, completing the proof. ∎
@@
Theorem 6.5 shows how the one-layer transformer model can interpolate $m$ input-output pairs through the following mechanism: (1) the embedding and self-attention sub-layers map distinct sequences to distinct vectors, (2) the hidden FNN layer maps these vectors to a larger space where they are linearly independent, (3) the outer FNN layer is solved as a system of linear equations.
@@
Theorem 6.5 also suggests an algorithm for finding an interpolating solution. First, note that every minor of $\psi(W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$ is real analytic on an open neighborhood of zero. Thus, it works to (1) sample $\theta$ from any continuous probability distribution, (2) choose $\epsilon>0$ sufficiently small so that the entries of $\epsilon W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top}$ are each in the interval of convergence of $\psi$, and (3) solve for $V$ in $Y=V^{\top}\psi(W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$.
@@
To compare the lower bound of Theorem 6.5 to the upper bound of Lemma 4.2, we have to assume a maximum context length $T$. Also, let $d_{r}=d_{1}\leavevmode\nobreak\ \forall r\in[m_{0}]$ for simplicity. Then, the total number of parameters is $\omega m+m(d+1)+2m_{0}(d_{0}+d_{1})d+(\omega+T)d$. Thus, the ratio between the upper and lower bounds is
$1+\frac{d+2}{\omega-1}+\frac{2m_{0}(d_{0}+d_{1})d}{(\omega-1)m}+\frac{(\omega+T)d}{(\omega-1)m},$ (7)
@@
which is $O(1)$ if and only if the number of parameters in the FNN is at least a constant proportion of the total number of parameters. To see that this is generally the case, consider how, in practice, $\min\{\omega,m\}\geq d\geq\max\{d_{0},d_{1}\}$ and $\omega m\geq Td$, so the ratio reduces to $O(m_{0}d^{2}/(\omega m))$, which is more clearly $O(1)$.
@@ 7 Experiment
We also provide an experiment showing that $\Omega(n)$ parameters seems to be sufficient not only in terms of memory capacity, but also in terms of training. In Figure 1, it can be seen that as the width of the hidden dimension (referred to as $m$ in the paper) of our decoder-only transformer model increases, the training error decreases. Furthermore, the number of parameters required to reach the entropy lower bound seems to be roughly linear, suggesting that first-order optimization methods can find an interpolating solution almost as soon as one exists.
@@
All experiments are performed on the TinyStories data set (Eldan and Li, 2023), where we take subsets of size 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000 from the data set. These data subsets have 377, 699, 952, 1223, 1459, 1637, 1896, 2117, 2374, 2637 unique contexts, respectively. Note that in our line plot, the minimum threshold varies for each data subset and is computed by taking the entropy lower bound and multiplying it by 0.01 (this helps to account for the fact that larger data sets have larger entropy lower bounds). We train 8 transformer models for each data subset, with $m$ values 4, 8, 16, 32, 64, 128, 256, 512. For simplicity and reduced computation time, we truncate each sequence to have a length of 10. Finally, we use an embedding size of 16 and the sinusoidal positional embedding from Vaswani et al (2017).
The implementation of our model is a decoder-only transformer and does not have any encoder blocks. It has only one decoder block, which consists of a self-attention layer followed by a linear layer, after which a GELU activation function is applied. Only one head is used for the self-attention. The output of this decoder block is fed into another linear layer, which produces the final output. We include a bias in both linear layers. Unlike many of the more complicated transformer models, we do not use any dropout, layer normalization, or residual connections in our implementation.
To compute the training error on each iteration, the cross-entropy loss function is used. For optimization, we use the Adam optimizer (Kingma and Ba, 2015) in full batch mode, with a stepsize of 0.0001 and no regularization (sometimes referred to as weight decay). Our training error values are computed based on 50,000 iterations. We choose a large number of iterations since our data subsets are small and we want to see how close we can get to the entropy lower bound for each value of $m$. Finally, we run our experiments on a computer with a Intel(R) Core(TM) i9-10900X CPU @ 3.70GHz processer and a NVIDIA RTX A4000 GPU with 16GB of memory. We refer you to the following Github link for the code to run our experiments: https://github.com/curtfox/decoder-memory-capacity.
@@ 8 Token-averaging
In the proof of Theorem 6.5, all that was required of the embedding and self-attention sub-layers was that the composition of them maps injectively into \$\0\mathbb{R}\backslash\{0\}$. This suggests a simple alternative to self-attention. Let $Z\in\mathbb{R}^{d\times\omega}$, $u\in\mathbb{R}^{\mathbb{N}}$, $W\in\mathbb{R}^{d\times m}$, $b\times\mathbb{R}^{m}$, and $V\in\mathbb{R}^{m\times\omega}$. Let $\theta$ denote $(Z,u,W,b,V)$. Let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at a point and not a polynomial there. Define
$h_{\theta}:\cup_{t=1}^{\infty}[\omega]^{t}$ $\to\mathrm{ri}\Delta^{\omega-1}$
@@
We call this a token-averaged FNN. Essentially, we have extended the FNN to be defined on $\cup_{t=1}^{\infty}[\omega]^{t}$ through the mapping $\alpha\mapsto Z[:,\alpha]u[1:|\alpha|]$, which we call token-averaging, and which we replaced self-attention with. It turns out this simple model also has optimal next-token prediction capacity, as we now prove.
Theorem 8.1.
Let $\omega,T\in\mathbb{N}$. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\omega}\times\mathbb{R}^{T}\times\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}$
@@
First, let $A$ denote $\cup_{t=1}^{T}[\omega]^{t}$ and define
@@
Note that the set of $(z,u)$ for which $f(z,u,\cdot)$ does not map injectively into \$\0\mathbb{R}\backslash\{0\}$ is precisely the zero set of $g$. Thus, the result follows if and only if $g$ is nontrivial if and only if (1) each $f(\cdot,\cdot,\alpha)$ is nontrivial and (2) each $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. Note that $f(\mathbbm{1}_{\omega},\mathbbm{1}_{T},\alpha)=|\alpha|$. This proves (1). It also proves (2) in the case where $|\alpha|\neq|\beta|$. All that is left is to prove (2) in the case where $|\alpha|=|\beta|$. Let $\alpha,\beta\in A$ such that $\alpha\neq\beta$ and $\tau\coloneqq|\alpha|=|\beta|$. Then there exists $s\in[\tau]$ such that $\alpha_{s}\neq\beta_{s}$. Choose such an $s$. Set $z=(1,\ldots,\omega)$ and $u=e_{s}$. Then $f(z,u,\alpha)=\alpha_{s}\neq\beta_{s}=f(z,u,\beta)$, proving (2), and so proving the theorem. ∎
Theorem 8.2.
Let $d,m,\omega\in\mathbb{N}$. For all $\theta\coloneqq(Z,u,W,b,V)\in\mathbb{R}^{d\times\omega}\times\mathbb{R}^{\mathbb{N}}\times\mathbb{R}^{d\times m}\times\mathbb{R}^{m}\times\mathbb{R}^{m\times\omega}$, define
$h_{\theta}:\cup_{t=1}^{\infty}[\omega]^{t}$ $\to\mathbb{R}^{\omega}$
@@
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\omega]^{t}$ be distinct and define $T=\max_{i\in[n]}|\alpha^{i}|$. Invoking Theorem 8.1, choose $(z,u)$ such that $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Invoking Theorem 6.4, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y^{\prime}\in(\mathrm{ri}\Delta^{\omega-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\omega\times n}$ such that $\phi(Y)=Y^{\prime}$ where $\phi$ is applied column-wise. Furthermore, there exists $V\in\mathbb{R}^{m\times\omega}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$. So, by defining $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d}$ and $W=\mathbbm{1}_{d}w^{\top}/\sqrt{d}$, we get that $Y=[h_{\theta}(\alpha^{i})]_{i\in[n]}$, completing the proof. ∎

### v3

@@ 6.3 Interpolation from the FNN
Combining Theorem 17 and Theorem 20 leads to a next-token prediction capacity lower bound for one-layer multi-head decoder-only transformers.
Theorem 21.
Let $d,m_{0},d_{0},m,\zeta\in\mathbb{N}$ and $d_{r}\in\mathbb{N}\leavevmode\nobreak\ \forall r\in[m_{0}]$. Let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Let $h_{\theta}$ be the one-layer multi-head decoder-only transformer model defined in Equation (LABEL:eq:model). Then the next-token prediction capacity of $\{h_{\theta}\mid\theta\}$ is greater than or equal to $m$.
@@
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\zeta]^{t}$ be distinct and define $T=\max_{i\in[n]}|\alpha^{i}|$. Invoking Theorem 17, choose $(z,u)$ such that $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Invoking Theorem 20, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y^{\prime}\in(\mathrm{ri}\Delta^{\zeta-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\zeta\times n}$ such that $\phi(Y)=Y^{\prime}$ where $\phi$ is the column-wise softmax function. Furthermore, there exists $V\in\mathbb{R}^{m\times\zeta}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$. So, by defining $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d},U=\mathbbm{1}_{d}u^{\top}/\sqrt{d},W_{1,r}=W_{2,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{r}}^{\top}/\sqrt{dd_{r}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{3,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{0}}^{\top}/\sqrt{dd_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}=\mathbbm{1}_{m_{0}d_{0}}\mathbbm{1}_{d}^{\top}/(m_{0}\sqrt{dd_{0}}),W=\mathbbm{1}_{d}w^{\top}/\sqrt{d},b=\eta\mathbbm{1}_{m}$, we get that $Y^{\prime}=[h_{\theta}(\alpha^{i})]_{i\in[n]}$, completing the proof. ∎
@@
Theorem 21 shows how the one-layer transformer model can interpolate $m$ input-output pairs through the following mechanism: (1) the embedding and self-attention sub-layers map distinct sequences to distinct vectors, (2) the hidden FNN layer maps these vectors to a larger space where they are linearly independent, (3) the outer FNN layer is solved as a system of linear equations.
@@
Theorem 21 also suggests an algorithm for finding an interpolating solution. First, note that every minor of $\psi(W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$ is real analytic on an open neighborhood of zero. Thus, it works to (1) sample $\theta$ from any continuous probability distribution, (2) choose $\epsilon>0$ sufficiently small so that the entries of $\epsilon W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top}$ are each in the interval of convergence of $\psi$, and (3) solve for $V$ in $Y=V^{\top}\psi(W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$.
Note that this is not a useful algorithm in the empirical setting because it requires that $\hat{p}$ be explicitly computed. And we are not claiming that we should try to use a transformer in the same way that it is used in our proof. In fact, as we will see in the next section, the embedding and self-attention sub-layers seem to be useful for more than just injectivity, which is, of course, expected. However, the proof does suggest a conceptual framework for guiding the design of network architectures. As long as each part of the proof continues to be satisfied, there is no reason not to swap out sub-layers for new mappings that have not been tried before. For example, in Section 8 we will show that the self-attention sub-layer can be swapped with something much simpler without loss of the next-token prediction capacity lower bound. While it may not be a practical substitution, it suggests that there may be more practical substitutions out there.
@@
Finally, it is time to circle back to Section 4 and compare the lower bound of Theorem 21 to the upper bound of Lemma 9. In order to do so, we have to assume a maximum context length $T$. Also, let $d_{r}=d_{1}\leavevmode\nobreak\ \forall r\in[m_{0}]$ for simplicity. Then, the total number of parameters is $\zeta m+m(d+1)+2m_{0}(d_{0}+d_{1})d+(\zeta+T)d$. Thus, the ratio between the upper and lower bounds is
$1+\frac{d+2}{\zeta-1}+\frac{2m_{0}(d_{0}+d_{1})d}{(\zeta-1)m}+\frac{(\zeta+T)d}{(\zeta-1)m},$ (5)
@@
which is $O(1)$ if and only if the number of parameters in the FNN is at least a constant proportion of the total number of parameters. To see that this is generally the case, consider how, in practice, $\min\{\zeta,m\}\geq d\geq\max\{d_{0},d_{1}\}$ and $\zeta m\geq Td$, so the ratio reduces to $O(m_{0}d^{2}/(\zeta m))$, which is more clearly $O(1)$.
@@ 7 Experiment
We also provide an experiment showing that $\Omega(n)$ parameters seems to be sufficient not only in terms of memory capacity, but also in terms of training. Taking a subset from a text corpus, we train the transformer model of Section 5 on it. We increase the hidden dimension, $m$, until we can get the training error close to the entropy lower bound. By increasing the subset size, we can see how the number of parameters required to reach the entropy lower bound depends on the number of unique contexts. In the line plot of Figure 1, the dependence seems to be roughly linear, suggesting that first-order optimization methods can find an interpolating solution almost as soon as one exists. Note that increasing the subset size not only increases the number of unique contexts, but also increases the vocabulary size, and so increases the number of parameters. Thus, $m$ does not have to increase very much to maintain the same training error, as can be seen in the heatmap of Figure 1. To see more clearly how $m$ affects the training error, we repeat the experiment but only train the FNN parameters this time. Again, we see a roughly linear dependence in the line plot of Figure 2, but we can also see that dependence more clearly in the heatmap. Note that training only the FNN parameters is similar to our proof technique where the embedding and self-attention sub-layers were only used for mapping injectively, not for training. The experiment shows that while this can work in practice, the slope in the line plot is roughly five times larger when only the FNN parameters are trained, so it is preferable to train all the parameters.
@@
Figure 2: Even when only training the FNN layers, we show that our model can memorize an increasing number of unique contexts as the hidden dimension $m$ is increased. Left: As the hidden dimension, $m$, increases and the number of unique contexts, $n$, decreases, the gap between the training error and the entropy lower bound trends downwards. Right: As the number of unique contexts increases, the minimum number of parameters required for the gap between the training error and the entropy lower-bound to fall below the minimum threshold increases (this trains both FNN linear layers, will check if only training the last layer works as well and let you know).
All experiments are performed on the TinyStories data set [56], where we take subsets of size 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000 from the data set and tokenize using WordPunctTokenizer from the Natural Language Toolkit. For simplicity and reduced computation time, we truncate each sequence to have a length of 10. These data subsets have 377, 699, 952, 1223, 1459, 1637, 1896, 2117, 2374, 2637 unique contexts, respectively. Finally, the corresponding vocabulary sizes are 182, 290, 358, 417, 469, 513, 560, 606, 648, 690, respectively. Note that in our line plot in Figure 1, the minimum threshold varies for each data subset and is computed by taking the entropy lower bound and multiplying it by 0.01 (this helps to account for the fact that larger data sets have larger entropy lower bounds). We perform the same calculation for the line plot in Figure 2 instead multiplying by 0.05. We train 8 transformer models for each data subset, with $m$ values 4, 8, 16, 32, 64, 128, 256, 512. Finally, we use an embedding size of 16 and the sinusoidal positional embedding from [5].
The implementation of our model is a decoder-only transformer and does not have any encoder blocks. It has only one decoder block, which consists of a self-attention layer followed by a linear layer, after which a GELU activation function is applied. Only one head is used for the self-attention. The output of this decoder block is fed into another linear layer, which produces the final output. We include a bias in both linear layers. Unlike many of the more complicated transformer models, we do not use any dropout, layer normalization, or residual connections in our implementation. The total number of parameters is $\zeta m+17(\zeta+m)+1088$ where $\zeta$ is the vocabulary size. The total number of FNN parameters is $\zeta m+\zeta+17m$. The proportion of parameters that are in the FNN is around two thirds for the data points in the line plot of Figure 1, and increases from around two thirds to 94% for the data points in the line plot of Figure 2.
To compute the training error on each iteration, the cross-entropy loss function is used. For optimization, we use the Adam optimizer [57] in full batch mode, with a stepsize of 0.0001 and no regularization (sometimes referred to as weight decay). Our training error values are computed based on 50,000 iterations. We choose a large number of iterations since our data subsets are small and we want to see how close we can get to the entropy lower bound for each value of $m$. Finally, we run our experiments on a computer with a Intel(R) Core(TM) i9-14900X processer and a NVIDIA GeForce RTX 4090 GPU with 24GB of memory. We refer you to the following Github link for the code to run our experiments: https://github.com/curtfox/decoder-memory-capacity.
@@ 8 Token-averaging
In the proof of Theorem 21, all that was required of the embedding and self-attention sub-layers was that the composition of them maps injectively into \$\0\mathbb{R}\backslash\{0\}$. This suggests a simple alternative to self-attention. Let $Z\in\mathbb{R}^{d\times\zeta}$, $u\in\mathbb{R}^{\mathbb{N}}$, $W\in\mathbb{R}^{d\times m}$, $b\times\mathbb{R}^{m}$, and $V\in\mathbb{R}^{m\times\zeta}$. Let $\theta$ denote $(Z,u,W,b,V)$. Let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at a point and not a polynomial there. Define
$h_{\theta}:\cup_{t=1}^{\infty}[\zeta]^{t}$ $\to\mathrm{ri}\Delta^{\zeta-1}$
@@
We call this a token-averaged FNN. Essentially, we have extended the FNN to be defined on $\cup_{t=1}^{\infty}[\zeta]^{t}$ through the mapping $\alpha\mapsto Z[:,\alpha]u[1:|\alpha|]$, which we call token-averaging, and which we replaced self-attention with. It turns out this simple model also has optimal next-token prediction capacity, as we now prove.
Theorem 22.
Let $\zeta,T\in\mathbb{N}$. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\zeta}\times\mathbb{R}^{T}\times\cup_{t=1}^{T}[\zeta]^{t}\to\mathbb{R}$
@@
First, let $A$ denote $\cup_{t=1}^{T}[\zeta]^{t}$ and define
@@
Note that the set of $(z,u)$ for which $f(z,u,\cdot)$ does not map injectively into \$\0\mathbb{R}\backslash\{0\}$ is precisely the zero set of $g$. Thus, the result follows if and only if $g$ is nontrivial if and only if (1) each $f(\cdot,\cdot,\alpha)$ is nontrivial and (2) each $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. Note that $f(\mathbbm{1}_{\zeta},\mathbbm{1}_{T},\alpha)=|\alpha|$. This proves (1). It also proves (2) in the case where $|\alpha|\neq|\beta|$. All that is left is to prove (2) in the case where $|\alpha|=|\beta|$. Let $\alpha,\beta\in A$ such that $\alpha\neq\beta$ and $\tau\coloneqq|\alpha|=|\beta|$. Then there exists $s\in[\tau]$ such that $\alpha_{s}\neq\beta_{s}$. Choose such an $s$. Set $z=(1,\ldots,\zeta)$ and $u=e_{s}$. Then $f(z,u,\alpha)=\alpha_{s}\neq\beta_{s}=f(z,u,\beta)$, proving (2), and so proving the theorem. ∎
Theorem 23.
Let $d,m,\zeta\in\mathbb{N}$. For all $\theta\coloneqq(Z,u,W,b,V)\in\mathbb{R}^{d\times\zeta}\times\mathbb{R}^{\mathbb{N}}\times\mathbb{R}^{d\times m}\times\mathbb{R}^{m}\times\mathbb{R}^{m\times\zeta}$, define
$h_{\theta}:\cup_{t=1}^{\infty}[\zeta]^{t}$ $\to\mathbb{R}^{\zeta}$
@@
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\zeta]^{t}$ be distinct and define $T=\max_{i\in[n]}|\alpha^{i}|$. Invoking Theorem 22, choose $(z,u)$ such that $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Invoking Theorem 20, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y^{\prime}\in(\mathrm{ri}\Delta^{\zeta-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\zeta\times n}$ such that $\phi(Y)=Y^{\prime}$ where $\phi$ is applied column-wise. Furthermore, there exists $V\in\mathbb{R}^{m\times\zeta}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$. So, by defining $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d}$ and $W=\mathbbm{1}_{d}w^{\top}/\sqrt{d}$, we get that $Y=[h_{\theta}(\alpha^{i})]_{i\in[n]}$, completing the proof. ∎